OpenStudy (anonymous):

2(m-8)=4(m+6) What is M

OpenStudy (anonymous):

@Lexi_Loves

OpenStudy (anonymous):

i would like to work it out so i don't have to always ask

OpenStudy (lexi_loves):

first you would have to do the stuff in ( )

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

then can't we change it to 2m - 2*8

OpenStudy (lexi_loves):

then we can change it into m-14 cuz, 2*8= 14 then we do the same thing to the other side!

OpenStudy (anonymous):

@Lexi_Loves

OpenStudy (anonymous):

hi my comp lags so hard XD

OpenStudy (anonymous):

so where does the two go does 2m =m?

OpenStudy (lexi_loves):

no ok. ... so this is your problem and you have to find M?! 2(m-8)=4(m+6) i just mulitmlied 2(8)=14 and 4(6)=24 2m-14=4m+24

OpenStudy (lexi_loves):

Do you know what to do now?

OpenStudy (anonymous):

no

OpenStudy (lexi_loves):

PEMDAS! do you know what that is?

OpenStudy (anonymous):

multiply 2m and 4m then

OpenStudy (lexi_loves):

i think your supposed to do the oppposite opperation to the 2m and 4m which would be what?

OpenStudy (anonymous):

divide

OpenStudy (lexi_loves):

good.

OpenStudy (anonymous):

im still kinda confused

OpenStudy (anonymous):

im still kinda confused

OpenStudy (lexi_loves):

on what?

OpenStudy (anonymous):

variables in expressions. so whats next

OpenStudy (lexi_loves):

ok I just Re-did my calculations and starting from the 2m-14=4m+24 you would subtract not multipli sorry. lol 2m-14=4m+24 all i did here was -14 from both sides 24-14=10 2m=4m+ 10

OpenStudy (lexi_loves):

no we have to get m by itself.

OpenStudy (anonymous):

ok

OpenStudy (lexi_loves):

and to do that we would have to divide by 4 on each side so, 2m=4m+10 /4 /4 cause we want the 10 to be alone as well.

OpenStudy (anonymous):

ok

OpenStudy (lexi_loves):

so what do you get?

OpenStudy (anonymous):

2m=m+10

OpenStudy (lexi_loves):

no if we were to divide 4 to both sides 2m=4m+10 all I did here is divide 4 by 2 4/2=.5 .5m=10

OpenStudy (anonymous):

ok

OpenStudy (lexi_loves):

now just do 10/.5 and see what you get!

OpenStudy (anonymous):

20

OpenStudy (lexi_loves):

GOOD JOB!! :D thats what i got!:D

OpenStudy (anonymous):

yay i aced the test thanks to you <3

OpenStudy (lexi_loves):

lol np tag me if you need anything at alll!!:)

OpenStudy (anonymous):

thanks alright bye X3